OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (2, 7) and (-6, -1)? A.(4,3) B.(-2,4) C.(-2,3)

OpenStudy (mathmate):

The centre is the mid-point of the line segment joining (2,7) and (-6,-1). Do you know how to find the midpoint of a line segment?

OpenStudy (anonymous):

no can you explain

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

The coordinates of the mid point between two segments can be found be the average of each of the x and y coordinates.

OpenStudy (anonymous):

so im thinking its 4,3

OpenStudy (mathmate):

How did you find (4,3)?

OpenStudy (anonymous):

idk i need help

OpenStudy (mathmate):

What is the x-coordinate of (2,7), and what is the y-coordinate?

OpenStudy (anonymous):

x=2 y=7

OpenStudy (mathmate):

How about (-6,-1)?

OpenStudy (anonymous):

x=-6 y=-1

OpenStudy (mathmate):

Good, so can you now find the average of x and y coordinates from these two points?

OpenStudy (anonymous):

-4,6

OpenStudy (mathmate):

The y-coordinate is correct. You may want to check on the x-coordinate.

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

no just 4

OpenStudy (mathmate):

one x=2 and the other x=-6. So the average is (2+ -6)/2 = (2-6)/2 = -4/2 = -2

OpenStudy (anonymous):

oh okay